### v1

# Sign changes of Kloosterman sums with moduli having at most two prime factors

Abstract.
We prove that the Kloosterman sum $\text{Kl}(1,q)$ changes sign infinitely many times, as $q\rightarrow+\infty$ with at most two prime factors. As a consequence, our result is unconditional compared with Drappeau and Maynard’s (Proc. Amer. Math. Soc., 2019), in which the existence of Laudau-Siegel zeros is required. Our arguments contain the Selberg sieve method, spectral theory and distribution of Kloosterman sums along with previous works by Fouvry, Matomäki, Michel, Sivak-Fischler, and Xi.
Keywords Kloosterman sum; Sign change; Selberg sieve method; Spectral theory; Equidistribution
MSC(2020) 11L05, 11N36(11N75, 11L20, 26D15)
1. Introduction
1.1. Background
Let $q$ be a fixed positive integer. For arbitrary integers $m$ and $n$, the classical Kloosterman sums are defined by
$S(m,n;q)=\sum_{a(q)\atop(a,q)=1}e\left(\frac{ma+n\overline{a}}{q}\right),$
where $e(x)=e(2\pi ix/q)$. Kloosterman sums have originated from Poincaré [12] and Kloosterman [8]. Scholars are concerned about whether there exists some specific explicit expression or asymptotic formula? Unfortunately, these results are beyond our capabilities for now.
Kloosterman sums are one of the central research interests in analytic number theory and related to numerous applications in Diophantine equations and automorphic forms. A well-known estimate from Weil [16] for the individual Kloosterman sum is
$\left|S(m,n;p)\right|\leq 2p^{1/2}(m,n,p)^{1/2},$
in general
$\left|S(m,n;q)\right|\leq q^{1/2}(m,n,q)^{1/2}\tau(q),$
and the factor $\tau(q)$ above can be modified to $2^{\omega(q)}$($\omega(q)$ represents the number of different prime factors of $q$) by Estermann [2].
The equidistributions of Kloosterman sums are proposed by Katz [6] and analogized by the Sato-Tate conjecture of elliptic curves, which we state as
Conjecture 1.1.
For each $f\in\mathcal{C}([0,\pi])$ and non-zero integer $a$, we have
$\lim_{x\rightarrow+\infty}\frac{1}{\pi(x)}\sum_{p\leq x}f\left(\theta_{p}(a)\right)=\frac{2}{\pi}\int_{0}^{\pi}f(\theta)\sin^{2}{\theta}d\theta,$
where $\theta_{p}(a)\in[0,\pi]$ is so-called Kloosterman sum angle defined as
$p^{1/2}\text{Kl}(a;p):=S(a,1;p)=2p^{1/2}\cos{\theta_{p}(a)}.$
Later, we usually refer to the above conjecture as the “horizontal” Sato-Tate conjecture. Katz inferred that the angles $\theta_{p}(a)$(for each $a$) are equidistributed, as $x\rightarrow+\infty$, with respect to the Sato-Tate measure
$\mu_{\text{ST}}=\frac{2}{\pi}\sin^{2}{\theta}d\theta.$
In [7], Katz showed that the numbers
$\left\{\theta_{p}(a):1\leq a<p\right\}$
equidistribute with respect to the same measure, as $p\rightarrow+\infty$, and we call it the “vertical” Sato-Tate law. Furthermore, a direct corollary of the conjecture is that Kloosterman sums change signs infinitely many times as $p\rightarrow+\infty$.
In this paper, we mainly concern about the sign changes Kloosterman sums. Fouvry and Michel [3][4] pioneering proved that
|$\left|\left\{X<q\leq 2X:\text{Kl}(1;q)\gtrless 0,\omega(q)\leq 23,q\ \text{square-free}\right\}\right|\gg\frac{X}{\log X}.$
Subsequently, this result was improved in a whole series of papers by Sivak-Fischler [14, 15], Matomäki [10], Xi [17, 18, 19] by replacing 23 with 18, 15, 10 and 7.
Recently, Drappeau and Maynard [1] proved that under the condition of the existence of Landau-Siegel zero, 23 can be reduced to 2. Motivated by their previous works, we showed that 23 can be improved to 2, unconditionally.
Our new approach is somewhat similar to Maynard’s idea [11] of applying a variant of the Selberg sieve to improve Zhang’s breakthrough [20] in the Twin Prime Conjecture without changing the level of the Bombieri-Vinogradov Theorem.
1.2. Our result
Our result is stated as the following.
Theorem 1.2.
Let $q$ be any square-free number with $\omega(q)\leq 2$. Then $\text{Kl}(1;q)$ changes sign infinitely many times, as $q\rightarrow+\infty$.
Notations Before we start, we need to give some notations:
$\bullet$ $U\ll V$ or $U=O(V)$, means $|U|\leq cV$ for some constant $c>0$.
$\bullet$ If not specified, we use $a\ (q)$ to represent $a(\bmod\ q)$ and $\overline{a}$ to represent the multiplicative inverse element modulo $q$.
$\bullet$ $\mathcal{L}=\log X$.
$\bullet \mathbb{C}$ represents the complex plane.
$\bullet$ $\Re s$ represents the real part of $s$.
$\bullet$ $\binom{a}{b}$ is binomial coefficient.
$\bullet$ $\int_{(c)}$ represents $\int_{c-i\infty}^{c+i\infty}$.
2. Preliminary
2.1. A new sieve weight
Define Selberg sieve weight $\left(\lambda_{d}\right)$ and state as
$\begin{cases}\lambda_{1}=1,\\ \left|\lambda_{d}\right|\leq 1,\\ \lambda_{d}=0,\ \text{if}\ d>\sqrt{D},\end{cases}$
where $D$ is the level of Selberg sieve. More specifically, we set $D=X^{1/2-\epsilon}$, and
$\lambda_{d}=\mu(d)F\left(\frac{\log\left(\sqrt{D}/d\right)}{\log\left(\sqrt{D}\right)}\right),$
where $F$ is a fixed smooth function supported on $[0,1]$ and vanishes at 0 to a suitable order (related to $\kappa$ and $l$ below).
Our purpose is to show
|$\left|\left\{X<q\leq 2X:\text{Kl}(1;q)\gtrless 0,\omega(q)\leq 2,q\ \text{square-free}\right\}\right|\geq c_{0}\frac{X}{\log X},$
where $c_{0}>0$ is a constant. It suffices to prove
$R^{\pm}(X):=\sum_{n}g\left(\frac{n}{X}\right)\mu^{2}(n)\frac{\left|S(1,1;n)\right|\pm S(1,1;n)}{\sqrt{n}}\left\{\rho-\left(\frac{\kappa}{2}\right)^{\omega(n)}\right\}\left(\sum_{d\mid n\Pi_{l}}\lambda_{d}\right)^{2}>0,$
where $g(X)$ is a fixed smooth function supported on [1,2], $\Pi_{l}$ is the product of the first $l$ primes.
In this paper, we specifically take
(1) $\kappa=4,\ l=10,\ F(x)=x^{\kappa+l}.$
Comparing with [17] and [18], we replaced $d\mid n$ with $d\mid n\Pi_{l}$. This idea is essentially due to Selberg [13], and our purpose of doing this is to more freely select the coefficients of $R_{1}(X)$ and $R_{3}(X)$ in the following argument.
In reviewing Selberg’s method within the context of the twin prime conjecture, as outlined in [13], we encounter a sieve weight given by
$\left(\sum_{d\mid n(n+2)}\lambda_{d}\right)^{2}.$
However, in the specific problem we are addressing, the prime factorization of $n+2$ remains indeterminate. Consequently, we aim to substitute $n+2$ with a finite product of primes. Furthermore, examining the proof presented in Section 5 of [18], we observe that in the sieve weight, the oscillation amplitude of $\lambda_{d}$ becomes significant when $d$ has small prime factors. To handle this technical issue, we introduce $\Pi_{l}$ to amplify the magnitude of possible cancellation coming from $\lambda_{d}$.
Alternatively, if we persist with the original sieve weight summed over divisors $d$, potential improvements to our results might be achieved by improving the level of the Selberg sieve, particularly in the context of Lemma 3.2. Nevertheless, enhancing this level presents a formidable challenge for us at present. Fortunately, by adopting another new sieve weight, we can circumvent this challenge.
2.2. Outline of the proof
Now we hope $R^{\pm}(X)$ has a lower bound larger than 0, as shown below
$R^{\pm}(X)\geq\rho R_{1}(X)\pm\rho R_{2}(X)-2R_{3}(X)>0,$
where
$R_{1}(X):=\sum_{n}g\left(\frac{n}{X}\right)\mu^{2}(n)\left|\text{Kl}(1;n)\right|\left(\sum_{d\mid n\Pi_{l}}\lambda_{d}\right)^{2},$
$R_{2}(X):=\sum_{n}g\left(\frac{n}{X}\right)\mu^{2}(n)\text{Kl}(1;n)\left(\sum_{d\mid n\Pi_{l}}\lambda_{d}\right)^{2},$
$R_{3}(X):=\sum_{n}g\left(\frac{n}{X}\right)\mu^{2}(n)\left|\text{Kl}(1;n)\right|\left(\frac{\kappa}{2}\right)^{\omega(n)}\left(\sum_{d\mid n\Pi_{l}}\lambda_{d}\right)^{2}.$
We will discuss $R_{1}(X)$, $R_{2}(X)$ and $R_{3}(X)$ in Section 4, 5 and Section 6, respectively. The relevant conclusion of $R_{1}(X)$ can be referred to [10] and [17]. $R_{2}(X)$ can be derived from a B-V Theorem for Kloosterman sum, which can be found in [4], [15] and [17]. For $R_{3}(X)$, it can be handled in a similar way as in [18].
To prove the Theorem 1.2, we need the following conclusions.
Proposition 2.1.
For any sufficiently large number $X$, there exists some constant $C_{1}>0$ such that
$R_{1}(X)\geq(1+o(1))C_{1}\frac{X}{\left(\log X\right)^{21}}.$
Specifically, we choose
$C_{1}=\left(4^{10}10!\binom{14}{10}\prod_{p\mid\Pi_{10}}\log p\right)^{2}\times 0.0142\tilde{g}(1),$
where $g(X)$ is a fixed smooth function defined above.
Proposition 2.2.
For any sufficiently large number $X$, we have
$R_{2}(X)=o\left(\frac{X}{\left(\log X\right)^{21}}\right).$
Proposition 2.3.
For any sufficiently large number $X$, there exists some constant $C_{2}>0$ such that
$R_{3}(X)\leq C_{2}\frac{X}{\left(\log X\right)^{21}}(1+o(1)).$
Specifically, we choose
$C_{2}=4^{21}2^{10}\left(\prod_{p\mid\Pi_{10}}\log p\right)^{2}\times 8817.853\tilde{g}(1).$
Proposition 2.1, 2.2, 2.3 implies Theorem 1.2 immediately by choosing a suitable $\rho$ (for instance $\rho=5$, since $C_{1}>2C_{2}$) such that $52^{\omega(n)}<$ 5, which leads to $\omega(n)\leq 2$.
3. Some Lemmas
The following lemmas are needed.
Lemma 3.1.
For a sufficiently large real number $X$, define
$H(X)=\sum_{n}g\left(\frac{n}{X}\right)\mu^{2}(n)\frac{\left|S(1,1;n)\right|}{\sqrt{n}}\left(\sum_{d\mid n}\lambda_{d}\right)^{2}.$
Thus we have
$H(X)\geq\tilde{g}(1)\sum_{2\leq i\leq 5}2^{i}A_{i}(F)C_{i}\cdot\frac{X}{\log X}(1+o(1)),$
where $C_{2}=0.11109$, $C_{3}=0.03557$, $C_{4}=0.01184$ and $C_{5}=0.00396$,
$A_{i}(F)=\mathop{\int\dots\int}_{R_{i}}\frac{L_{i}^{2}\left(F;X^{1-\alpha_{2}-\dots-\alpha_{i}},X^{\alpha_{2}},\dots,X^{\alpha_{i}}\right)}{\alpha_{2}\dots\alpha_{i}(1-\alpha_{2}-\dots-\alpha_{i})}\text{d}\alpha_{2}\dots\text{d}\alpha_{i},$
$L_{i}(F;\alpha_{1},\alpha_{2},\dots,\alpha_{i})=\sum_{\mathcal{A}\subseteq\left\{\alpha_{1},\alpha_{2},\dots,\alpha_{i}\right\}\atop\sum_{\alpha\in\mathcal{A}}\alpha<\frac{1}{4}}(-1)^{|\mathcal{A}|}F\left(1-4\sum_{\alpha\in\mathcal{A}}\alpha\right),$
and
$R_{2}:=$$\left\{\alpha_{2}\in[\eta,1):\left(\frac{3}{4}+\eta\right)(1-\alpha_{2})<\alpha_{2}<\frac{1}{2}\right\},$
$R_{3}:=$$\left\{(\alpha_{2},\alpha_{3})\in[\eta,1)^{2}:\frac{1}{2}(1-\alpha_{2}-\alpha_{3})<\alpha_{2},\ \alpha_{3}<\alpha_{2}<1-\alpha_{2}-\alpha_{3}\right\},$
$R_{4}:=$$\left\{(\alpha_{2},\alpha_{3},\alpha_{4})\in[\eta,1)^{3}:\frac{1}{2}(1-\alpha_{2}-\alpha_{3}-\alpha_{4})<\alpha_{2}+\alpha_{3}\right\}$
$\cap\left\{(\alpha_{2},\alpha_{3},\alpha_{4})\in[\eta,1)^{3}:\alpha_{4}<\alpha_{3}<\alpha_{2}<1-\alpha_{2}-\alpha_{3}-\alpha_{4}\right\},$
$R_{5}:=$$\left\{(\alpha_{2},\alpha_{3},\alpha_{4},\alpha_{5})\in[\eta,1)^{4}:\frac{1}{2}(1-\alpha_{2}-\alpha_{3}-\alpha_{4}-\alpha_{5})<\alpha_{2}+\alpha_{3}+\alpha_{4}\right\}$
$\cap\left\{(\alpha_{2},\alpha_{3},\alpha_{4},\alpha_{5})\in[\eta,1)^{4}:\frac{1}{2}(\alpha_{3}+\alpha_{4}+\alpha_{5})<\alpha_{2}\right\}$
$\cap\left\{(\alpha_{2},\alpha_{3},\alpha_{4},\alpha_{5})\in[\eta,1)^{4}:\alpha_{5}<\alpha_{4}<\alpha_{3}<\alpha_{2}<1-\alpha_{2}-\alpha_{3}-\alpha_{4}\alpha_{5}\right\},$
$\eta:=$ $10^{-2023}.$
Proof.
See Proposition 2.1 in [10] or Proposition 2.1 in [17]. ∎
Lemma 3.2.
For any $A>0$, there exists some $B=B(A)>0$ such that
$\sum_{q\leq\sqrt{X}\mathcal{L}^{-B}}3^{\omega(q)}\left|\sum_{n\equiv 0(q)}\mu^{2}(n)g\left(\frac{n}{X}\right)\text{Kl}(1;n)\right|\ll X\mathcal{L}^{-A},$
where the implied constant depends on $A$ and $g$.
Proof.
See Lemma 2.1 in [17]. ∎
Before presenting the following lemmas, let’s first establish some definitions. Define a smooth function $F(x)$ and assume it has a Taylor expansion
$F(x):=\sum_{n\geq 0}\frac{a_{n}}{n!}x^{n}$
with $F(0)=0$. For $N>1$ and $\0s\in\mathbb{C}\backslash\left\{0\right\}$, we further define
$\check{F}_{N}(s):=\sum_{n\geq 0}\frac{a_{n}}{(s\log N)^{n}}.$
Lemma 3.3.
Let $N>1$ be not an integer. For any coefficient $y_{n}$ with $\ll\tau(n)^{O(1)}(\log n)^{O(1)}$, we have
$\sum_{n\leq N}y_{n}F\left(\frac{\log(N/n)}{\log N}\right)=\frac{1}{2\pi i}\int_{2-i\infty}^{2+i\infty}\check{F}_{N}(s)Y(s)\frac{N^{s}}{s}\text{d}s,$
where
$Y(s)=\sum_{n\geq 0}\frac{y_{n}}{n^{s}},\ \Re s>1.$
Proof.
See Lemma 2.1 in [9].∎
Lemma 3.4.
Let $v$, $v_{1}$, $v_{2}$ be fixed positive integers, $m$ be fixed even positive integer, $M>1$ be any real number. Suppose $P$, $Q$ are two smooth functions that have zeros of orders at least $v_{1}+m/2$, $v_{2}+m/2$ at 0, respectively. Let $Z(s_{1},s_{2})$ be holomorphic in the right half plane containing a neighborhood of $(0,0)$ with
$\left(\frac{\text{d}^{n}}{\text{d}s^{n}}Z(s,s\xi)\right)_{s=0}=0,\ 0\leq n<m,$
and
$\left(\frac{\text{d}^{m}}{\text{d}s^{m}}Z(s,s\xi)\right)_{s=0}=C_{0}\xi^{\frac{m}{2}}Z^{(m)}(0,0)\neq 0,$
where $|\xi|=2$ is a circle in $\mathbb{C}$, $C_{0}$ is a constant. Put
$R:=\mathop{\text{Res}}_{{(s_{1},s_{2})=(0,0)}}\check{P}_{M}(s_{1})\check{Q}_{M}(s_{2})Z(s_{1},s_{2})s_{1}^{v_{1}-1}s_{2}^{v_{2}-1}\frac{M^{s_{1}+s_{2}}}{(s_{1}+s_{2})^{v}}.$
Thus we have
$R=(1+o(1))Z^{(m)}(0,0)\frac{(\log M)^{v-v_{1}-v_{2}-m}}{\Gamma(v)m!}\int_{0}^{1}P^{(v_{1}+m/2)}(x)Q^{(v_{2}+m/2)}(x)(1-x)^{v-1}\text{d}x.$
Proof.
We follow the proof of Lemma A.2 in [18] and use Goldston et al.’s method in [5]. Assume
$P(x):=\sum_{k\geq 0}\frac{a_{k}}{k!}x^{k},\ Q(x):=\sum_{k\geq 0}\frac{b_{k}}{k!}x^{k}.$
For $M>1$ and $\0s\in\mathbb{C}\backslash\left\{0\right\}$, define
$\check{P}_{M}(s)=\sum_{k\geq 0}\frac{a_{k}}{(s\log M)^{k}},\ \check{Q}_{M}(s)=\sum_{k\geq 0}\frac{b_{k}}{(s\log M)^{k}}.$
Therefore, we have
(2) $R=\sum_{k_{1}\geq v_{1}}\sum_{k_{2}\geq v_{2}}\frac{a_{k_{1}}b_{k_{2}}}{(\log M)^{k_{1}+k_{2}}}\mathop{\text{Res}}_{{(s_{1},s_{2})=(0,0)}}Z(s_{1},s_{2})\frac{M^{s_{1}+s_{2}}}{s_{1}^{k_{1}-v_{1}+1}s_{2}^{k_{2}-v_{2}+1}(s_{1}+s_{2})^{v}},$
and the residue therein is equal to
$\frac{1}{(2\pi i)^{2}}\int_{C_{2}}\int_{C_{1}}\frac{Z(s_{1},s_{2})M^{s_{1}+s_{2}}}{s_{1}^{k_{1}-v_{1}+1}s_{2}^{k_{2}-v_{2}+1}(s_{1}+s_{2})^{v}}\text{d}s_{1}\text{d}s_{2},$
where $C_{1}$, $C_{2}$ are the circles $|s_{1}|=\rho$ and $|s_{2}|=2\rho$ with a small $\rho>0$. We write $s_{1}=s$, $s_{2}=s\xi$, then the double integral can be rewritten as
(3) $\frac{1}{(2\pi i)^{2}}\int_{C_{3}}\int_{C_{1}}\frac{Z(s,s\xi)M^{s(1+\xi)}}{s^{v+k_{1}-v_{1}+k_{2}-v_{2}+1}\xi^{k_{2}-v_{2}+1}(\xi+1)^{v}}\text{d}s\text{d}\xi,$
where $C_{3}$ is the circle $|\xi|=2$.
For $s$, we find
$\frac{1}{(v+k_{1}-v_{1}+k_{2}-v_{2})!}\left(\frac{\text{d}^{v+k_{1}-v_{1}+k_{2}-v_{2}}}{\text{d}s^{v+k_{1}-v_{1}+k_{2}-v_{2}}}Z(s,s\xi)M^{s(1+\xi)}\right)_{s=0}.$
From the condition of Lemma 3.4 this can be expanded as
$\sum_{i=m}^{v+k_{1}-v_{1}+k_{2}-v_{2}}\binom{v+k_{1}-v_{1}+k_{2}-v_{2}}{i}$ $\left(\frac{\text{d}^{i}}{\text{d}s^{i}}Z(s,s\xi)\right)_{s=0}$
×$\times(\xi+1)^{v+k_{1}-v_{1}+k_{2}-v_{2}-i}(\log M)^{v+k_{1}-v_{1}+k_{2}-v_{2}-i}.$
We only need to consider the term with the quantity $(\log M)^{v+k_{1}-v_{1}+k_{2}-v_{2}-m}$, thus substituting into (3) to derive
$\frac{1}{(v+k_{1}-v_{1}+k_{2}-v_{2})!}\binom{v+k_{1}-v_{1}+k_{2}-v_{2}}{m}$ $Z^{(m)}(0,0)(\log M)^{v+k_{1}-v_{1}+k_{2}-v_{2}-m}$
(4) ×$\times\frac{1}{2\pi i}\int_{C_{3}}\frac{(\xi+1)^{k_{1}-v_{1}+k_{2}-v_{2}-m}}{\xi^{k_{2}-v_{2}-m/2+1}}\text{d}\xi.$
For the integral in (3), we will deal with the following four cases as
$\begin{cases}&k_{1}-v_{1}\geq\frac{m}{2}\\ &k_{2}-v_{2}\geq\frac{m}{2},\end{cases}\ \begin{cases}&k_{1}-v_{1}<\frac{m}{2}\\ &k_{2}-v_{2}<\frac{m}{2},\end{cases}\ \begin{cases}&k_{2}-v_{2}\geq\frac{m}{2}\\ &k_{1}-v_{1}<\frac{m}{2},\end{cases}\ \begin{cases}&k_{2}-v_{2}>\frac{m}{2}\\ &k_{1}-v_{1}\geq\frac{m}{2}.\end{cases}$
The integral values for the last three cases are all equal to 0. When $k_{1}-v_{1}<\frac{m}{2}$ and $k_{2}-v_{2}<\frac{m}{2}$, only one pole $\xi=-1$ of order $m-(k_{1}-v_{1})-(k_{2}-v_{2})$ is in the region bounded by the circle $|\xi|=2$. Then Cauchy’s residue theorem implies that
$\frac{1}{2\pi i}\int_{C_{3}}\frac{\xi^{m/2-(k_{2}-v_{2})-1}}{(\xi+1)^{m-(k_{1}-v_{1})-(k_{2}-v_{2})}}\text{d}\xi=\mathop{\text{Res}}_{\xi=-1}\frac{\xi^{m/2-(k_{2}-v_{2})-1}}{(\xi+1)^{m-(k_{1}-v_{1})-(k_{2}-v_{2})}}=0.$
When $k_{1}-v_{1}<\frac{m}{2}$ and $k_{2}-v_{2}\geq\frac{m}{2}$, we divide this case into two cases, namely
$\begin{cases}&k_{2}\geq v_{2}+\frac{m}{2}+\left(v_{1}+\frac{m}{2}-k_{1}\right)\\ &k_{1}<v_{1}+\frac{m}{2},\end{cases}\ \begin{cases}&v_{2}+\frac{m}{2}\leq k_{2}<v_{2}+\frac{m}{2}+\left(v_{1}+\frac{m}{2}-k_{1}\right)\\ &k_{1}<v_{1}+\frac{m}{2}.\end{cases}$
For the first one, we see the integral is equal to
$\frac{1}{2\pi i}\int_{C_{3}}\frac{(\xi+1)^{k_{1}+k_{2}-v_{1}-v_{2}-m}}{\xi^{k_{2}-v_{2}-m/2+1}}\text{d}\xi.$
Only one pole $\xi=0$ of order $k_{2}-v_{2}-m/2+1$ is in the region bounded by $|\xi|=2$. Thus Cauchy’s residue theorem implies that
$\frac{1}{2\pi i}\int_{C_{3}}\frac{(\xi+1)^{k_{1}+k_{2}-v_{1}-v_{2}-m}}{\xi^{k_{2}-v_{2}-m/2+1}}\text{d}\xi=\mathop{\text{Res}}_{\xi=0}\frac{(\xi+1)^{k_{1}+k_{2}-v_{1}-v_{2}-m}}{\xi^{k_{2}-v_{2}-m/2+1}}=0.$
For the second one, the integral is
$\frac{1}{2\pi i}\int_{C_{3}}\frac{1}{(\xi+1)^{v_{1}+v_{2}+m-k_{1}-k_{2}}\xi^{k_{2}-v_{2}-m/2+1}}\text{d}\xi,$
and two poles $\xi=-1$ and $\xi=0$ are in the region bounded by $|\xi|=2$. Surprisingly, in this case, we can obtain
$\mathop{\text{Res}}_{\xi=-1}\frac{1}{(\xi+1)^{v_{1}+v_{2}+m-k_{1}-k_{2}}\xi^{k_{2}-v_{2}-m/2+1}}=-\mathop{\text{Res}}_{\xi=0}\frac{1}{(\xi+1)^{v_{1}+v_{2}+m-k_{1}-k_{2}}\xi^{k_{2}-v_{2}-m/2+1}},$
thus the integral is still equal to 0.
The last case can be handled in a similar way to the third one. For the convenience of calculation, in this case, the positions of $s_{1}$ and $s_{2}$ are better to be interchanged, i.e., let $|s_{1}|=2\rho$, $|s_{2}|=\rho$.
Combining the above, only the case
$\begin{cases}&k_{1}-v_{1}\geq\frac{m}{2}\\ &k_{2}-v_{2}\geq\frac{m}{2}\end{cases}$
remains to be dealt with. The integral contributes $\binom{k_{1}-v_{1}+k_{2}-v_{2}-m}{k_{2}-v_{2}-m/2}$ and makes (2)
$(1+o(1))Z^{(m)}(0,0)\frac{(\log M)^{v-v_{1}-v_{2}-m}}{m!}$
×$\ \ \ \ \ \ \ \times\sum_{k_{1}\geq v_{1}+m/2}\sum_{k_{2}\geq v_{2}+m/2}\frac{a_{k_{1}}b_{k_{2}}}{(v+k_{1}-v_{1}+k_{2}-v_{2}-m)!}\binom{k_{1}-v_{1}+k_{2}-v_{2}-m}{k_{2}-v_{2}-m/2}$
$=$ $(1+o(1))Z^{(m)}(0,0)\frac{(\log M)^{v-v_{1}-v_{2}-m}}{m!\Gamma(v)}\int_{0}^{1}P^{(v_{1}+m/2)}(x)Q^{(v_{2}+m/2)}(x)(1-x)^{v-1}\text{d}x,$
where we have used Lemma A.2 (Page 1225, line 3) in [18] for the last step. ∎
4. Lower bound for $R_{1}(X)$
For the convenience of calculation, we only consider the sum restricted to the set as we mentioned in Lemma 3.1 which is
$P_{2}(X):=\left\{p_{1}p_{2}\sim X|X^{\eta}<p_{1}<p_{2},\ p_{1}>p_{2}^{3/4}X^{\eta}\right\}.$
This set corresponds to $R_{2}$ in the lemma 3.1 and $\eta$ is the same as in Lemma 3.1. Recall the definition of $\lambda_{d}$, in this case we see
$p_{1},p_{2}>\sqrt{D}.$
Therefore, a lower bound of $R_{1}(X)$ can be obatined, which is
$R_{1}(X)=$ $\sum_{n}g\left(\frac{n}{X}\right)\mu^{2}(n)\left|\text{Kl}(1;n)\right|\left(\sum_{d\mid n\Pi_{l}}\lambda_{d}\right)^{2}$
$\geq$ $\sum_{n\in P_{2}(X)}g\left(\frac{n}{X}\right)\mu^{2}(n)\left|\text{Kl}(1;n)\right|\left(\sum_{d\mid n\Pi_{l}}\lambda_{d}\right)^{2}$
$=$ $\left(\sum_{d\mid\Pi_{l}}\lambda_{d}\right)^{2}\sum_{n\in P_{2}(X)}g\left(\frac{n}{X}\right)\mu^{2}(n)\left|\text{Kl}(1;n)\right|$
$\geq$ $(1+o(1))\left(\sum_{d\mid\Pi_{l}}\lambda_{d}\right)^{2}\times 4\tilde{g}(1)A_{2}(F)C_{2}\cdot\frac{X}{\log X}.$
In view of the selection of (1), we find that all $1/(\log X)^{j}$ vanishes in
$\left(\sum_{d\mid\Pi_{l}}\lambda_{d}\right)^{2},$
unless $j\geq 10$. For example, the term related to $1/\log X$ is equal to
$\frac{\binom{14}{1}}{\log\sqrt{D}}$ $\left(\log 2+\log 3+\dots+\log 29\right.$
$\left.-\log(2\cdot 3)-\log(2\cdot 29)-\dots-\log(23\cdot 29)-\dots-\log(2\cdot 3\cdot\dots\cdot 29)\right)=0.$
For the terms related to $1/(\log X)^{10}$, only
$10!\binom{14}{10}\left(\prod_{p\mid\Pi_{10}}\log p\right)\cdot\frac{1}{(\log\sqrt{D})^{10}}$
remains. Thus, we obtain the lower bound is
$(1+o(1))\left(\frac{10!\binom{14}{10}\prod_{p\mid\Pi_{10}}\log p}{(\log\sqrt{D})^{10}}\right)^{2}\times 4\tilde{g}(1)A_{2}(F)C_{2}\cdot\frac{X}{\log X}$
$=$ $(1+o(1))\left(4^{10}10!\binom{14}{10}\prod_{p\mid\Pi_{10}}\log p\right)^{2}\times 0.0142\tilde{g}(1)\frac{X}{(\log X)^{21}},$
where the numerical value of $A_{2}(F)=0.0319586...$ can be found in [[17], Proposition 2.1] or [[18], Section 4].
5. Upper bound for $R_{2}(X)$
We write $R_{2}(X)$ as
$R_{2}(X)=$ $\sum_{n}g\left(\frac{n}{X}\right)\mu^{2}(n)\text{Kl}(1;n)\left(\sum_{d\mid n\Pi_{l}}\lambda_{d}\right)^{2}$
$=$ $\sum_{n}g\left(\frac{n}{X}\right)\mu^{2}(n)\text{Kl}(1;n)\sum_{d\mid n\Pi_{l}}\xi_{d}$
$=$ $\sum_{d\leq D}\xi_{d}\sum_{n\Pi_{l}\equiv 0(d)}g\left(\frac{n}{X}\right)\mu^{2}(n)\text{Kl}(1;n),$
where
$\xi_{d}=\sum_{d=[h_{1},h_{2}]}\lambda_{h_{1}}\lambda_{h_{2}},\ |\xi_{d}|\leq 3^{\omega(d)}.$
The definition of $\xi_{d}$ indicates that $d$ is a square-free number. Let $d=d_{1}d_{2}$ satisfying $(d,\Pi_{l})=d_{1}$, we also have $(d_{2},\Pi_{l}/d_{1})=1$. Then $R_{2}(X)$ is
$\sum_{d_{1}\mid\Pi_{l}}\sum_{d_{2}\leq D/d_{1}\atop(d_{2},\Pi_{l}/d_{1})=1}\xi_{d_{1}d_{2}}\sum_{n\equiv 0(d_{2})}g\left(\frac{n}{X}\right)\mu^{2}(n)\text{Kl}(1;n)$
$\ll$ $\sum_{d_{1}\mid\Pi_{l}}3^{\omega(d_{1})}\sum_{d_{2}}3^{\omega(d_{2})}\left|\sum_{n\equiv 0(d_{2})}g\left(\frac{n}{X}\right)\mu^{2}(n)\text{Kl}(1;n)\right|$
$\ll$ $\frac{X}{(\log X)^{A}}=o\left(\frac{X}{(\log X)^{21}}\right),$
where we adopt the Lemma 3.2. Thus Proposition 2.1 is proved.
6. Upper bound for $R_{3}(X)$
The proof of Proposition 2.3 is actually similar to Section 5 in [18]. However, for completeness, we still provide the details of the proof.
6.1. Separation of variables and establishment of integrals
Recall the Weil’s bound for Kloosterman sums, we get an upper bound of $R_{3}(X)$ which is
$\sum_{n}g\left(\frac{n}{X}\right)\mu^{2}(n)\kappa^{\omega(n)}\left(\sum_{d\mid n\Pi_{l}}\lambda_{d}\right)^{2}$
$=$ $\sum_{\Pi_{l}=P_{0}P_{1}}\sum_{n\equiv 0(P_{0})\atop(n/P_{0},P_{1})=1}g\left(\frac{n}{X}\right)\mu^{2}(n)\kappa^{\omega(n)}\left(\sum_{d\mid n\Pi_{l}}\lambda_{d}\right)^{2}$
$=$ $\sum_{P_{0}\mid\Pi_{l}}\mu^{2}(P_{0})\kappa^{\omega(P_{0})}\sum_{(n,\Pi_{l})=1}g\left(\frac{nP_{0}}{X}\right)\mu^{2}(n)\kappa^{\omega(n)}\left(\sum_{d\mid n\Pi_{l}}\lambda_{d}\right)^{2}$
$=$ $\sum_{P_{0}\mid\Pi_{l}}\mu^{2}(P_{0})\kappa^{\omega(P_{0})}\sum_{d}\xi_{d}\sum_{n\Pi_{l}\equiv 0(d)\atop(n,\Pi_{l})=1}g\left(\frac{nP_{0}}{X}\right)\mu^{2}(n)\kappa^{\omega(n)}$
$=$ $\sum_{P_{0}\mid\Pi_{l}}\mu^{2}(P_{0})\kappa^{\omega(P_{0})}\sum_{d}\xi_{d}\sum_{dk\equiv 0(\Pi_{l})\atop(d,k)=1,(dk/\Pi_{l},\Pi_{l})=1}g\left(\frac{dkP_{0}}{X\Pi_{l}}\right)\mu^{2}\left(\frac{dk}{\Pi_{l}}\right)\kappa^{\omega\left(\frac{dk}{\Pi_{l}}\right)}$
$=$ $\sum_{P_{0}\mid\Pi_{l}}\mu^{2}(P_{0})\kappa^{\omega(P_{0})}\sum_{\Pi_{l}=\Delta_{1}\Delta_{2}}\sum_{d\equiv 0(\Delta_{1})\atop(d/\Delta_{1},\Pi_{l})=1}\xi_{d}$
(5) ×$\ \ \ \ \ \ \ \ \ \ \ \ \ \ \times\sum_{k\equiv 0(\Delta_{2})\atop(k,d)=1,(k/\Delta_{2},\Pi_{l})=1}g\left(\frac{dkP_{0}}{X\Delta_{1}\Delta_{2}}\right)\mu^{2}\left(\frac{dk}{\Delta_{1}\Delta_{2}}\right)\kappa^{\omega\left(\frac{dk}{\Delta_{1}\Delta_{2}}\right)}.$
We state the inner sum in (6.1) as
$\sum_{(k,\Pi_{l})=1,(k,d)=1}g\left(\frac{(d/\Delta_{1})kP_{0}}{X}\right)\mu^{2}\left(\frac{d}{\Delta_{1}}k\right)\kappa^{\omega\left(dk/\Delta_{1}\right)}$
$=$ $\frac{1}{2\pi i}\int_{(2)}\tilde{g}(s)\left(\sum_{(k,(d/\Delta_{1})\Pi_{l})=1}\frac{\mu^{2}(dk/\Delta_{1})\kappa^{\omega(dk/\Delta_{1})}}{(d/\Delta_{1})^{s}k^{s}P_{0}^{s}}\right)X^{s}\text{d}s,$
where the Mellin inverse transform of $\tilde{g}$ is used. Substituting into (6.1) yields
$\frac{X}{2\pi i}\int_{(1)}\tilde{g}(s+1)\left(\sum_{P_{0}\mid\Pi_{l}}\frac{\mu^{2}(P_{0})\kappa^{\omega(P_{0})}}{P_{0}^{s+1}}\right)$ $\left(\sum_{\Pi_{l}=\Delta_{1}\Delta_{2}}\sum_{d\equiv 0(\Delta_{1})\atop(d/\Delta_{1},\Pi_{l})=1}\frac{\xi_{d}\mu^{2}(d/\Delta_{1})\kappa^{\omega(d/\Delta_{1})}}{(d/\Delta_{1})^{s+1}}\right)$
(6) ×$\times\left(\sum_{(k,d\Pi_{l}/\Delta_{1})=1}\frac{\mu^{2}(k)\kappa^{\omega(k)}}{k^{s+1}}\right)X^{s}\text{d}s.$
Furthermore, using the Euler product formula for the Dirichlet series in the third bracket we derive
$\sum_{(k,d\Pi_{l}/\Delta_{1})=1}\frac{\mu^{2}(k)\kappa^{\omega(k)}}{k^{s+1}}$
$=$ $\prod_{p\nmid d\Pi_{l}/\Delta_{1}}\left(1+\frac{\kappa}{p^{s+1}}\right)$
$=$ $\prod_{p}\left(1+\frac{\kappa}{p^{s+1}}\right)\left(1-\frac{1}{p^{s+1}}\right)^{\kappa}\zeta^{\kappa}(s+1)\prod_{p\mid d\Pi_{l}/\Delta_{1}}\left(1+\frac{\kappa}{p^{s+1}}\right)^{-1}.$
Let
$G(s+1):=\prod_{p}\left(1+\frac{\kappa}{p^{s+1}}\right)\left(1-\frac{1}{p^{s+1}}\right)^{\kappa}.$
The sum in the first bracket of (6.1) is equal to
$\prod_{p\mid\Pi_{l}}\left(1+\frac{\kappa}{p^{s+1}}\right),$
thus we have
(7) $\frac{X}{2\pi i}\int_{(1)}\tilde{g}(s+1)G(s+1)\zeta^{\kappa}(s+1)N(s)X^{s}\text{d}s,$
where
$N(s):=\sum_{\Pi_{l}=\Delta_{1}\Delta_{2}}\sum_{d\equiv 0(\Delta_{1})\atop(d/\Delta_{1},\Pi_{l})=1}\frac{\xi_{d}\mu^{2}(d/\Delta_{1})\kappa^{\omega(d/\Delta_{1})}}{(d/\Delta_{1})^{s+1}}\prod_{p\mid d/\Delta_{1}}\left(1+\frac{\kappa}{p^{s+1}}\right)^{-1}.$
6.2. Analysis of $N(s)$
Obviously, there is
$N(s)=\sum_{\Pi_{l}=\Delta_{1}\Delta_{2}}\sum_{(d,\Pi_{l})=1}\frac{\xi_{d\Delta_{1}}\mu^{2}(d)\kappa^{\omega(d)}}{d^{s+1}}\prod_{p\mid d}\left(1+\frac{\kappa}{p^{s+1}}\right)^{-1}.$
We define
$\beta(d,s):=\frac{d^{s+1}}{\mu^{2}(d)\kappa^{\omega(d)}}\prod_{p\mid d}\left(1+\frac{\kappa}{p^{s+1}}\right),$
then from the definition of $\xi_{d}$ we have
$N(s)=$ $\sum_{\Pi_{l}=\Delta_{1}\Delta_{2}}\sum_{(d,\Pi_{l})=1}\frac{\xi_{d\Delta_{1}}}{\beta(d,s)}$
$=$ $\sum_{\Pi_{l}=\Delta_{1}\Delta_{2}}\sum_{(d,\Pi_{l})=1}\left(\sum_{d\Delta_{1}=[d_{1},d_{2}]}\frac{\lambda_{d_{1}}\lambda_{d_{2}}}{\beta(d,s)}\right)$
$=$ $\sum_{\Pi_{l}=\Delta_{1}\Delta_{2}}\sum_{[d_{1},d_{2}]\equiv 0(\Delta_{1})\atop([d_{1},d_{2}]/\Delta_{1},\Pi_{l})=1}\frac{\lambda_{d_{1}}\lambda_{d_{2}}}{\beta([d_{1},d_{2}]/\Delta_{1},s)}.$
Since $d_{i}(i=1,2)$ are square-free integers, let $(d_{1},d_{2})=m$, $d_{1}=d_{1}^{\prime}m$ and $d_{2}=d_{2}^{\prime}m$. Then we know $m$, $d_{1}^{\prime}$ and $d_{2}^{\prime}$ are pairwise coprime (for convenience, we will use the symbol $d_{1}$, $d_{2}$ in the following text). Now we can write
$N(s)=\sum_{\Delta_{1}\mid\Pi_{l}}\sum_{m\leq\sqrt{D}}\sum_{md_{1}d_{2}\equiv 0(\Delta_{1})\atop(d_{1},d_{2})=1,(md_{1}d_{2}/\Delta_{1},\Pi_{l})=1}\frac{\lambda_{md_{1}}\lambda_{md_{2}}}{\beta(md_{1}d_{2}/\Delta_{1},s)}.$
It is beneficial to separate $\Delta_{1}$ into $\Delta_{1}^{\prime}\Delta_{2}^{\prime}\Delta_{3}$ satisfying $\Delta_{1}^{\prime}\mid m$, $\Delta_{2}^{\prime}\mid d_{1}$ and $\Delta_{3}\mid d_{2}$ (in the following text, we still use the symbol $\Delta_{1}$ and $\Delta_{2}$), which leads to
$N(s)=$ $\sum_{\Delta_{1}\Delta_{2}\Delta_{3}\mid\Pi_{l}}\sum_{m\leq\sqrt{D}}\sum_{m\equiv 0(\Delta_{1})\atop{d_{1}\equiv 0(\Delta_{2})\atop{d_{2}\equiv 0(\Delta_{3})\atop(d_{1},d_{2})=1,\left(\frac{m}{\Delta_{1}}\frac{d_{1}}{\Delta_{2}}\frac{d_{2}}{\Delta_{3}},\Pi_{l}\right)=1}}}\frac{\lambda_{md_{1}}\lambda_{md_{2}}}{\beta\left(\frac{m}{\Delta_{1}}\frac{d_{1}}{\Delta_{2}}\frac{d_{2}}{\Delta_{3}},s\right)}$
(8) $=$ $\sum_{\Delta_{1}\Delta_{2}\Delta_{3}\mid\Pi_{l}}\sum_{m\leq\sqrt{D}\atop{m\equiv 0(\Delta_{1})\atop(m/\Delta_{1},\Pi_{l})=1}}\frac{1}{\beta(m/\Delta_{1},s)}\sum_{d_{1}\equiv 0(\Delta_{2})\atop(d_{1}/\Delta_{2},\Pi_{l})=1}\frac{\lambda_{md_{1}}}{\beta(d_{1}/\Delta_{2},s)}\sum_{d_{2}\equiv 0(\Delta_{3})\atop{(d_{1},d_{2})=1\atop(d_{2}/\Delta_{3},\Pi_{l})=1}}\frac{\lambda_{md_{2}}}{\beta(d_{2}/\Delta_{3},s)}.$
Firstly, we deal with the inner sum in (6.2). From the definition of $\lambda_{d}$ we have
$\sum_{d_{2}\equiv 0(\Delta_{3})\atop{(d_{1},d_{2})=1\atop(d_{2}/\Delta_{3},\Pi_{l})=1}}\frac{\lambda_{md_{2}}}{\beta(d_{2}/\Delta_{3},s)}$
$=$ $\sum_{(d_{2},\Pi_{l})=1\atop(d_{1},d_{2})=1}\frac{\lambda_{md_{2}\Delta_{3}}}{\beta(d_{2},s)}$
$=$ $\mu(m\Delta_{3})\sum_{\left(d_{2},\frac{m}{\Delta_{1}}\frac{d_{1}}{\Delta_{2}}\Pi_{l}\right)=1}\frac{\mu(d_{2})}{\beta(d_{2},s)}F\left(\frac{\log(\sqrt{D}/md_{2}\Delta_{3})}{\log\sqrt{D}}\right).$
Set
$P(x):=F\left(x\cdot\frac{\log(\sqrt{D}/m\Delta_{3})}{\log\sqrt{D}}\right),$
and apply Lemma 3.3 to $P(x)$ (or see Page 1218 in [18]), we obtain
$\frac{\mu(m\Delta_{3})}{2\pi i}\int_{(2)}\check{P}_{\sqrt{D}/m\Delta_{3}}(t)\left(\sum_{\left(d_{2},\frac{m}{\Delta_{1}}\frac{d_{1}}{\Delta_{2}}\Pi_{l}\right)=1}\frac{\mu(d_{2})}{\beta(d_{2},s)d_{2}^{t}}\right)\frac{(\sqrt{D}/(m\Delta_{3}))^{t}}{t}\text{d}t.$
Note that
$\check{P}_{\sqrt{D}/m\Delta_{3}}(t)=\check{F}_{\sqrt{D}}(t),$
thus we have
$\frac{\mu(m\Delta_{3})}{2\pi i}\int_{(2)}\check{F}_{\sqrt{D}}(t)\left(\sum_{\left(d_{2},\frac{m}{\Delta_{1}}\frac{d_{1}}{\Delta_{2}}\Pi_{l}\right)=1}\frac{\mu(d_{2})}{\beta(d_{2},s)d_{2}^{t}}\right)\frac{(\sqrt{D}/(m\Delta_{3}))^{t}}{t}\text{d}t,$
and we state the sum in the integrand as
$\prod_{p\nmid\frac{m}{\Delta_{1}}\frac{d_{1}}{\Delta_{2}}\Pi_{l}}\left(1-\frac{1}{\beta(p,s)p^{t}}\right)$
$=$ $\prod_{p}\left(1-\frac{1}{\beta(p,s)p^{t}}\right)\prod_{p\mid\frac{m}{\Delta_{1}}\frac{d_{1}}{\Delta_{2}}\Pi_{l}}\left(1-\frac{1}{\beta(p,s)p^{t}}\right)^{-1}$
$=$ $\prod_{p\nmid\Pi_{l}}\left(1-\frac{1}{\beta(p,s)p^{t}}\right)\prod_{p\mid\frac{m}{\Delta_{1}}\frac{d_{1}}{\Delta_{2}}}\left(1-\frac{1}{\beta(p,s)p^{t}}\right)^{-1}$
(9) $=$ $\prod_{p\nmid\Pi_{l}}\left(1-\frac{1}{\beta(p,s)p^{t}}\right)\prod_{p\mid(m/\Delta_{1})}\left(1-\frac{1}{\beta(p,s)p^{t}}\right)^{-1}\prod_{p\mid(d_{1}/\Delta_{2})}\left(1-\frac{1}{\beta(p,s)p^{t}}\right)^{-1}.$
Substituting into (6.2) to derive
$\frac{1}{2\pi i}\int_{(2)}$ $\check{F}_{\sqrt{D}}(t_{1})\frac{\sqrt{D}^{t_{1}}}{t_{1}}\left(\sum_{\Delta_{1}\Delta_{2}\Delta_{3}\mid\Pi_{l}}\frac{\mu(\Delta_{3})}{\Delta_{3}^{t_{1}}}\prod_{p\nmid\Pi_{l}}\left(1-\frac{1}{\beta(p,s)p^{t_{1}}}\right)\right)$
$\times$ $\left(\sum_{m\leq\sqrt{D}\atop{m\equiv 0(\Delta_{1})\atop(m/\Delta_{1},\Pi_{l})=1}}\frac{\mu(m)}{\beta(m/\Delta_{1},s)m^{t_{1}}}\prod_{p\mid(m/\Delta_{1})}\left(1-\frac{1}{\beta(p,s)p^{t_{1}}}\right)^{-1}\right)$
(10) $\times$ $\left(\sum_{d_{1}\equiv 0(\Delta_{2})\atop(d_{1}/\Delta_{2},\Pi_{l})=1}\frac{\lambda_{md_{1}}}{\beta(d_{1}/\Delta_{2},s)}\prod_{p\mid(d_{1}/\Delta_{2})}\left(1-\frac{1}{\beta(p,s)p^{t_{1}}}\right)^{-1}\right)\text{d}t_{1}.$
Next, following much the same way, we deal with the inner sum in (6.2), which leads to
$\frac{\mu(m\Delta_{2})}{2\pi i}\int_{(2)}\check{F}_{\sqrt{D}}(t_{2})\left(\sum_{\left(d_{1},\frac{m}{\Delta_{1}}\Pi_{l}\right)=1}\frac{\mu(d_{1})}{\beta(d_{1},s)d_{1}^{t_{2}}}\prod_{p\mid d_{1}}\left(1-\frac{1}{\beta(p,s)p^{t_{1}}}\right)^{-1}\right)\frac{(\sqrt{D}/(m\Delta_{2}))^{t_{2}}}{t_{2}}\text{d}t_{2}.$
The sum in the integrand can be transformed to
$\prod_{p\nmid\Pi_{l}}\left(1-\frac{1}{\beta(p,s)p^{t_{2}}\left(1-\frac{1}{\beta(p,s)p^{t_{1}}}\right)}\right)\prod_{p\mid(m/\Delta_{1})}\left(1-\frac{1}{\beta(p,s)p^{t_{2}}\left(1-\frac{1}{\beta(p,s)p^{t_{1}}}\right)}\right)^{-1}.$
Thus we obtain that (6.2) is further equal to
$\frac{1}{(2\pi i)^{2}}\iint$ $\check{F}_{\sqrt{D}}(t_{1})\check{F}_{\sqrt{D}}(t_{2})\frac{\sqrt{D}^{t_{1}+t_{2}}}{t_{1}+t_{2}}$
×$\times\left(\sum_{\Delta_{1}\Delta_{2}\Delta_{3}\mid\Pi_{l}}\frac{\mu(\Delta_{2})\mu(\Delta_{3})}{\Delta_{3}^{t_{1}}\Delta_{2}^{t_{2}}}\prod_{p\nmid\Pi_{l}}\left(1-\frac{1}{\beta(p,s)p^{t_{1}}}-\frac{1}{\beta(p,s)p^{t_{2}}}\right)\right)$
×$\times\left(\sum_{m\leq\sqrt{D}\atop{m\equiv 0(\Delta_{1})\atop(m/\Delta_{1},\Pi_{l})=1}}\frac{\mu^{2}(m)}{\beta(m/\Delta_{1},s)m^{t_{1}+t_{2}}}\prod_{p\mid(m/\Delta_{1})}\left(1-\frac{1}{\beta(p,s)p^{t_{1}}}-\frac{1}{\beta(p,s)p^{t_{2}}}\right)^{-1}\right)\text{d}t_{1}\text{d}t_{2}$
$=\frac{1}{(2\pi i)^{3}}\iiint$ $\check{F}_{\sqrt{D}}(t_{1})\check{F}_{\sqrt{D}}(t_{2})\frac{\sqrt{D}^{t_{1}+t_{2}+w}}{t_{1}+t_{2}+w}$
×$\times\left(\sum_{\Delta_{1}\Delta_{2}\Delta_{3}\mid\Pi_{l}}\frac{\mu(\Delta_{2})\mu(\Delta_{3})}{\Delta_{3}^{t_{1}}\Delta_{2}^{t_{2}}}\prod_{p\nmid\Pi_{l}}\left(1-\frac{1}{\beta(p,s)p^{t_{1}}}-\frac{1}{\beta(p,s)p^{t_{2}}}\right)\right)$
×$\times\left(\sum_{m\equiv 0(\Delta_{1})\atop(m/\Delta_{1},\Pi_{l})=1}\frac{\mu^{2}(m)}{\beta(m/\Delta_{1},s)m^{t_{1}+t_{2}+w}}\prod_{p\mid(m/\Delta_{1})}\left(1-\frac{1}{\beta(p,s)p^{t_{1}}}-\frac{1}{\beta(p,s)p^{t_{2}}}\right)^{-1}\right)\text{d}t_{1}\text{d}t_{2}\text{d}w,$
where the third integral comes from the Mellin inverse transform of $1_{m\leq\sqrt{D}}$.
We rewrite the summation over $m$ as
$\frac{\mu^{2}(\Pi_{1})}{\Delta_{1}^{t_{1}+t_{2}+w}}\prod_{p\nmid\Pi_{l}}\left(1+\frac{1}{\beta(p,s)p^{t_{1}+t_{2}+w}}\left(1-\frac{1}{\beta(p,s)p^{t_{1}}}-\frac{1}{\beta(p,s)p^{t_{2}}}\right)^{-1}\right),$
which leads to
$N(s)=\frac{1}{(2\pi i)^{3}}\iiint$ $\check{F}_{\sqrt{D}}(t_{1})\check{F}_{\sqrt{D}}(t_{2})\frac{\sqrt{D}^{t_{1}+t_{2}+w}}{t_{1}+t_{2}+w}$
×$\times\left(\sum_{\Delta_{1}\Delta_{2}\Delta_{3}\mid\Pi_{l}}\frac{\mu^{2}(\Delta_{1})\mu(\Delta_{2})\mu(\Delta_{3})}{\Delta_{1}^{t_{1}+t_{2}+w}\Delta_{2}^{t_{2}}\Delta_{3}^{t_{1}}}\right)$
×$\times\prod_{p\nmid\Pi_{l}}\left(1-\frac{1}{\beta(p,s)p^{t_{1}}}-\frac{1}{\beta(p,s)p^{t_{2}}}+\frac{1}{\beta(p,s)p^{t_{1}+t_{2}+w}}\right)\text{d}t_{1}\text{d}t_{2}\text{d}w.$
Further, we can write
$\sum_{\Delta_{1}\Delta_{2}\Delta_{3}\mid\Pi_{l}}\frac{\mu^{2}(\Delta_{1})\mu(\Delta_{2})\mu(\Delta_{3})}{\Delta_{1}^{t_{1}+t_{2}+w}\Delta_{2}^{t_{2}}\Delta_{3}^{t_{1}}}$
$=$ $\sum_{\Delta_{1}b\mid\Pi_{l}}\frac{\mu(b)\mu^{2}(\Delta_{1})}{b^{t_{1}}\Delta_{1}^{t_{1}+t_{2}+w}}\sum_{\Delta_{2}\mid b}\frac{\mu^{2}(\Delta_{2})}{\Delta_{2}^{t_{2}-t_{1}}}$
$=$ $\sum_{\Delta_{1}b\mid\Pi_{l}}\frac{\mu(b)\mu^{2}(\Delta_{1})}{b^{t_{1}}\Delta_{1}^{t_{1}+t_{2}+w}}\prod_{p\mid b}\left(1+\frac{1}{p^{t_{2}-t_{1}}}\right)$
$=$ $\sum_{a\mid\Pi_{l}}\frac{\mu(a)}{a^{t_{1}}}\prod_{p\mid a}\left(1+\frac{1}{p^{t_{2}-t_{1}}}\right)\sum_{\Delta_{1}\mid a}\frac{\mu(\Delta_{1})}{\Delta_{1}^{t_{2}+w}}\prod_{p\mid\Delta_{1}}\left(1+\frac{1}{p^{t_{2}-t_{1}}}\right)^{-1}$
$=$ $\sum_{a\mid\Pi_{l}}\frac{\mu(a)}{a^{t_{1}}}\prod_{p\mid a}\left(1+\frac{1}{p^{t_{2}-t_{1}}}-\frac{1}{p^{t_{2}+w}}\right)$
$=$ $\prod_{p\mid\Pi_{l}}\left(1-\frac{1}{p^{t_{1}}}-\frac{1}{p^{t_{2}}}+\frac{1}{p^{t_{1}+t_{2}+w}}\right),$
and finally, we get
$N(s)=\frac{1}{(2\pi i)^{3}}\iiint$ $\check{F}_{\sqrt{D}}(t_{1})\check{F}_{\sqrt{D}}(t_{2})\frac{\sqrt{D}^{t_{1}+t_{2}+w}}{t_{1}+t_{2}+w}$
×$\times\prod_{p\mid\Pi_{l}}\left(1-\frac{1}{p^{t_{1}}}-\frac{1}{p^{t_{2}}}+\frac{1}{p^{t_{1}+t_{2}+w}}\right)$
×$\times\prod_{p\nmid\Pi_{l}}\left(1-\frac{1}{\beta(p,s)p^{t_{1}}}-\frac{1}{\beta(p,s)p^{t_{2}}}+\frac{1}{\beta(p,s)p^{t_{1}+t_{2}+w}}\right)\text{d}t_{1}\text{d}t_{2}\text{d}w.$
6.3. Upper bound for $R_{3}(X)$
From (7) and $N(s)$ we obtained above, there is
$\sum_{n}g\left(\frac{n}{X}\right)\mu^{2}(n)\kappa^{\omega(n)}\left(\sum_{d\mid n\Pi_{l}}\lambda_{d}\right)^{2}$
::=:= $\frac{X}{(2\pi i)^{4}}\iiiint\tilde{g}(s+1)G(s+1)\zeta^{\kappa}(s+1)\check{F}_{\sqrt{D}}(t_{1})\check{F}_{\sqrt{D}}(t_{2})$
(12) ×$\ \ \ \ \ \ \times H(s,t_{1},t_{2},w)I(s,t_{1},t_{2},w)\frac{X^{s}\sqrt{D}^{t_{1}+t_{2}+w}}{t_{1}t_{2}w}\text{d}s\text{d}t_{1}\text{d}t_{2}\text{d}w,$
where each integral is over $1+it$, $t\in\mathbb{R}$ and
$H(s,t_{1},t_{2},w):=\prod_{p}\left(1-\frac{1}{\beta(p,s)p^{t_{1}}}-\frac{1}{\beta(p,s)p^{t_{2}}}+\frac{1}{\beta(p,s)p^{t_{1}+t_{2}+w}}\right)\zeta^{\kappa}(s+1),$
$I(s,t_{1},t_{2},w):=$ $\prod_{p\mid\Pi_{l}}\left(1-\frac{1}{\beta(p,s)p^{t_{1}}}-\frac{1}{\beta(p,s)p^{t_{2}}}+\frac{1}{\beta(p,s)p^{t_{1}+t_{2}+w}}\right)^{-1}$
×$\times\prod_{p\mid\Pi_{l}}\left(1-\frac{1}{p^{t_{1}}}-\frac{1}{p^{t_{2}}}+\frac{1}{p^{t_{1}+t_{2}+w}}\right)$
for $\Re(s+t_{1}+t_{2}+w)>0$, $\Re(s+t_{1})>0$, $\Re(s+t_{2})>0$ and $\Re s>-1$.
Let
$K(s,t_{1},t_{2},w):=$ $H(s,t_{1},t_{2},w)$
×$\times\prod_{p}\left(1-\frac{1}{p^{s+t_{1}+1}}\right)^{-\kappa}\prod_{p}\left(1-\frac{1}{p^{s+t_{2}+1}}\right)^{-\kappa}\prod_{p}\left(1-\frac{1}{p^{s+t_{1}+t_{2}+w+1}}\right)^{\kappa}$
(13) ×$\times\left(\frac{\zeta(s+t_{1}+t_{2}+w+1)}{\zeta(s+t_{1}+1)\zeta(s+t_{2}+1)}\right)^{\kappa}\left(\frac{s(s+t_{1}+t_{2}+w)}{(s+t_{1})(s+t_{2})}\right)^{\kappa},$
then we have
$\sum_{n}g\left(\frac{n}{X}\right)\mu^{2}(n)\kappa^{\omega(n)}\left(\sum_{d\mid n\Pi_{l}}\lambda_{d}\right)^{2}$
$=$ $\frac{X}{(2\pi i)^{4}}\iiiint\tilde{g}(s+1)G(s+1)\zeta^{\kappa}(s+1)\check{F}_{\sqrt{D}}(t_{1})\check{F}_{\sqrt{D}}(t_{2})I(s,t_{1},t_{2},w)$
(14) ×$\ \ \ \ \ \ \times K(s,t_{1},t_{2},w)\left(\frac{(s+t_{1})(s+t_{2})}{s(s+t_{1}+t_{2}+w)}\right)^{\kappa}\frac{X^{s}\sqrt{D}^{t_{1}+t_{2}+w}}{t_{1}t_{2}w}\text{d}s\text{d}t_{1}\text{d}t_{2}\text{d}w.$
We will evaluate the multiple-integral by shifting contours. It can be checked that the quadruple integral appeared in (6.3) and Section 5 in [18] (Page 1220) are the same, except for a finite term $I(s,t_{1},t_{2},w)$. Therefore, after shifting the $w-$contour and $s-$contour through the same way, we only need to handle the integrand with $t_{1}$, $t_{2}$ as
$\tilde{g}(1)$ $G(1)\sum_{j=1}^{\kappa}\binom{\kappa}{j}\sum_{i=0}^{j-1}\binom{j-1}{i}(\log X)^{j-i-1}(-1)^{i}\frac{\Gamma(i+j)}{\Gamma^{2}(j)}$
×$\times\check{F}_{\sqrt{D}}(t_{1})\check{F}_{\sqrt{D}}(t_{2})K(0,t_{1},t_{2},0)I(0,t_{1},t_{2},0)(t_{1}t_{2})^{j-1}\frac{\sqrt{D}^{t_{1}+t_{2}}}{(t_{1}+t_{2})^{i+j}}.$
Note that
$K(0,0,0,0)=\frac{1}{G(1)},$
and $I(0,t_{1},t_{2},0)$ satisfies the condition in Lemma 3.4. Then from Lemma 3.4 with $m=2l$ we have
$\sum_{n}g\left(\frac{n}{X}\right)\mu^{2}(n)\kappa^{\omega(n)}\left(\sum_{d\mid n\Pi_{l}}\lambda_{d}\right)^{2}$
$=$ $(1+o(1))\frac{4^{2l+1}\tilde{g}(1)I^{(2l)}(0,0,0,0)}{(2l)!}\frac{X}{(\log X)^{2l+1}}\sum_{j=1}^{\kappa}\binom{\kappa}{j}\frac{1}{\Gamma^{2}(j)}$
×$\times\int_{0}^{1}F^{(j+l)}(x)^{2}(1-x)^{j-1}\left(\sum_{i=0}^{j-1}\binom{j-1}{i}(x-1)^{i}4^{j-1-i}\right)\text{d}x$
$=$ $(1+o(1))\frac{4^{2l+1}\tilde{g}(1)I^{(2l)}(0,0,0,0)}{(2l)!}\frac{X}{(\log X)^{2l+1}}$
×$\times\sum_{j=1}^{\kappa}\binom{\kappa}{j}\frac{1}{\Gamma^{2}(j)}\int_{0}^{1}F^{(j+l)}(x)^{2}(1-x)^{j-1}(x+3)^{j-1}\text{d}x.$
In view of the restrictions in (1) again, with the help of mathematical software, we find
$R_{3}(X)\leq(1+o(1))4^{21}2^{10}\left(\prod_{p\mid\Pi_{10}}\log p\right)^{2}\times 8817.853\tilde{g}(1)\frac{X}{(\log X)^{2l+1}},$
which claims the Proposition 2.3.
Acknowledgements
This work is supported by the National Natural Science Foundation of China (Nos. 12271320, 11871317), and the Natural Science Basic Research Plan for Distinguished Young Scholars in Shaanxi Province of China (No. 2021JC-29).